OpenStudy (anonymous):

What is the result of adding the system of equations? 2x + y = 4 3x - y = 6 x = 2 x = 10 5x = 10

OpenStudy (anonymous):

x = 2?

OpenStudy (just_one_last_goodbye):

give me a sec ok? ^_^

OpenStudy (just_one_last_goodbye):

Step: Solve2x+y=4for y: 2x+y+−2x=4+−2x(Add -2x to both sides) y=−2x+4 Step: Substitute−2x+4foryin3x−y=6: 3x−y=6 3x−(−2x+4)=6 5x−4=6(Simplify both sides of the equation) 5x−4+4=6+4(Add 4 to both sides) \[\frac{ 5x }{ 5 } = \frac{ 10 }{ 5 }\] (Divide both sides by 5) x=2 Step: Substitute2forxiny=−2x+4: y=−2x+4 y=(−2)(2)+4 y=0 (Simplify both sides of the equation)

OpenStudy (anonymous):

oh so it 5x=10?

OpenStudy (just_one_last_goodbye):

no :)

OpenStudy (just_one_last_goodbye):

5x/5 cancel out while dividing by 5

OpenStudy (just_one_last_goodbye):

y = 0 and x = 2

OpenStudy (anonymous):

oh ok thanks:)

OpenStudy (just_one_last_goodbye):

No prob ^_^